OpenStudy (anonymous):

How else can 5 and ^4/3 be written?

OpenStudy (anonymous):

do you mean \[5^{\frac{4}{3}}\] ?

OpenStudy (anonymous):

or do you mean \[5+\frac{4}{3}\]

OpenStudy (anonymous):

The first reply is correct.

OpenStudy (anonymous):

\[5^{\frac{4}{3}}=\sqrt[3]{5^4}\]